OpenStudy (anonymous):

What is the exact area of a circle whose diameter is 18 cm? A. 9 cm2 B. 18 cm2 C. 36 cm2 D. 81 cm2

OpenStudy (anonymous):

I think it is a, i will give medal @dan815

OpenStudy (anonymous):

@amistre64 @DolphinDawg @TheSmartOne @jdoe0001

OpenStudy (anonymous):

Diameter is 18 units Therefore radius r=9 units Area=Pi*r^2 =3.14*81 =254.34 sq units

OpenStudy (anonymous):

oh, i mean c

OpenStudy (anonymous):

i was thinking radius, whoops

TheSmartOne (thesmartone):

You copy-pasted your question and because of that you lost the pi in each answer because of copy-pasting errors.

OpenStudy (dan815):

haha my help is sold out for a medal!

TheSmartOne (thesmartone):

in each answer choice*

OpenStudy (anonymous):

yep, sorry bout that

OpenStudy (dan815):